divide and simplify
y^3 + 9y divided by y^2 + 2y - 35
----------------------------------
y^2 - 81 divided by y^2 - 14y + 45
I am not sure how to type this but it is suppose to look like a division problem with a line between in the middle.

Answer by london maths tutor(243)

[(y^3 + 9y) / (y^2 +2y - 35)] ÷ [(y^2-81) / (y^2 -14y + 45)]
Factorise various brackets:
y^3 + 9y = y(y^2 + 9)
y^2 +2y - 35 = (y+7)(y-5)
y^2-81 = (y+9)(y-9)
y^2 -14y + 45 = (y-9)(y-5)
Put the factorised brackets back to the question and change the question to multiplication:
[y(y^2+9)]/[(y+7)(y-5)] * [(y-9)(y-5)]/[(y+9)(y-9)]
cancel out "(y-9)" and "(y-5)"
This gives:
y(y^2+9)/[(y+7)(y+9)]